OpenStudy (anonymous):

Nina was curious about the height of the Eiffel Tower. She used a 1.2 m model of the tower and measured its shadow at 2 p.m. The length of the shadow was 0.9 m. She then measured the Eiffel Tower's shadow and found that it was 240 m. What is the height of the Eiffel Tower?

OpenStudy (anonymous):

SOMEBODY HELP ME

jimthompson5910 (jim_thompson5910):

solve for x 1.2/0.9 = x/240

jimthompson5910 (jim_thompson5910):

The fact that it was measured at 2:00 pm is irrelevant info (and not needed or used here)

OpenStudy (anonymous):

That is not an answer choice. the choices are a) 0.45 m b) 2,880 m c) 30 m d)320 m

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you have to solve to get your answer

jimthompson5910 (jim_thompson5910):

1.2/0.9 = x/240 1.2*240 = 0.9x ... ... ... x = ???

OpenStudy (anonymous):

oh yes its 320. thank you so much @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you got it